each leeg of an isoceles triangle exceed twice the base by 5 inches find each side of the triangle if the perimeter is 5 feet.
Answer by mananth(16946)

let base be x
leg = 2x+(5/12) feet
perimeter = 5 feet
..
x+2x+5/12+2x+5/12=5
5x+10/12 =5
5x= 5- 10/12
5x= (60-10) /12 ( LCD = 12)
5x=50/12
x= 50/5*12
x= 10/12
= 5/6 feet

leg = 2*5/6 +5/12
leg = 5/3 +5/12
leg = 20 +5 /12
leg =25/12